The giants left san francisco heading northeast at 210 km/hr. at the same time the patriots left boston and headed southwest traveling at 180km/hr. after 8 hours both teams landed at the airport in indianapolis at the same time, how far apart are boston and san francisco?
Answer by mananth(16949)

They are travelling in a straight line towards each other
combine their speeds
210+180=390 km/h
t=8
distance =d
d=r*t
d=390*8
d=3120 km